# On dense subalgebras of the singular ideal in groupoid C*-algebras

### v1

@@ 3. Groups with the Density Property
Let A be a $C^{*}$-algebra, $\mathcal{A}\subseteq A$ a dense -subalgebra, and $I\trianglelefteq A$ an ideal. Then $I\cap\mathcal{A}$ is dense in $I$ if and only if $I\cap\mathcal{A}$ contains a net $(u_{\beta})$ of contractions such that $\lim_{\beta}\|a-au_{\beta}\|=0$ for all $a\in I$. Moreover, $u_{\beta}$ can be chosen to be positive.
@@
Assume that $I\cap\mathcal{A}$ is dense in $I$. Since $I$ is a $C^{*}$-algebra, it contains an approximate unit $(u_{\beta})$. Then $I\cap\mathcal{A}$ is a dense -subalgebra, so each $u_{\beta}$ can be approximated by a sequence of positive elements in the unit ball of $I\cap\mathcal{A}$. By a diagonalisation argument, $I\cap\mathcal{A}$ contains a net of positive elements $I$ converging strongly to the identity in $I$. Conversely, assume that $I\cap\mathcal{A}$ contains a net of contractions converging strongly to the identity in $I$. Let $a\in I$ and $\varepsilon>0$. There exist $u\in I\cap\mathcal{A}$ and $b\in\mathcal{A}$ such that $\left\|u\right\|\leq 1$, $\left\|a-au\right\|<\frac{\varepsilon}{2}$ and $\left\|a-b\right\|<\frac{\varepsilon}{2}$. Then $bu\in I\cap\mathcal{A}$ and
We now prove our first application of this lemma.
Proposition 3.4.
Let $\Gamma$ be an amenable discrete group and suppose $\mathcal{X}_{1},\dots,\mathcal{X}_{n}\subseteq\mathrm{Sub}(\Gamma)$ are closed and conjugation-invariant sets of subgroups satisfying $\mathcal{X}_{i}\in\mathcal{D}_{\Gamma}$ for all $i=1,\dots,n$. Then, $\mathcal{X}\coloneq\bigcup_{i=1}^{n}\mathcal{X}_{i}$ satisfies $\mathcal{X}\in\mathcal{D}_{\Gamma}$.
@@
By definition, we have $J_{\Gamma,\mathcal{X}}=\bigcap^{n}_{i=1}J_{\Gamma,\mathcal{X}_{i}}$. For each $i=1,\dots,n$ there exists, by Lemma 3.3, a net $(u_{i,\beta})$ of positive contractions in $J_{\Gamma,\mathcal{X}_{i}}\cap\mathbb{C}[\Gamma]$ converging strongly to the identity in $J_{\Gamma,\mathcal{X}_{i}}$ and hence in $J_{\Gamma,\mathcal{X}}\subseteq J_{\Gamma,\mathcal{X}_{i}}$. Therefore, by a diagonalisation argument, the net $(u_{\beta_{1},\beta_{2}..,\beta_{n}}):=(u_{1,\beta_{1}}\cdot u_{2,\beta_{n}}\cdot...\cdot u_{n,\beta_{n}})$ (equipped with the lexicographic partial ordering on $(\beta_{1},...,\beta_{n})$) has a subnet $(u_{\gamma})$ converging strongly to the identity in $J_{\Gamma,\mathcal{X}}$. The net $(u_{\gamma})$ lies in $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$, so $\mathcal{X}\in\mathcal{D}_{\Gamma}$ by Lemma 3.3. ∎
Next, we show the Density Property is preserved under taking quotients of groups by normal subgroups. For an amenable discrete group $\Gamma$ and a normal subgroup $N$, let $Q_{N}:C^{*}_{r}(\Gamma)\to C^{*}_{r}(\Gamma/N)$ be the *-homomorphism induced by the quotient map $q_{N}:\Gamma\to\Gamma/N$. The map
$q^{*}_{N}:\text{Sub}(\Gamma/N)$ $\rightarrow\text{Sub}(\Gamma),\text{ }X\mapsto q_{N}^{-1}(X),$
is a continuous injection satisfying $\text{Conj}_{g}\circ q^{*}_{N}=q^{*}_{N}\circ\text{Conj}_{q_{N}(g)}$ for all $g\in\Gamma$, where $\text{Conj}_{g}$ denotes conjugation by $g$. Therefore, if $\mathcal{X}$ is a closed and conjugation-invariant set of subgroups of $\Gamma/N$, then $q^{*}_{N}(\mathcal{X})$ is a closed and conjugation-invariant set of subgroups of $\Gamma$.
Theorem 3.5.
Suppose $\Gamma$ is an amenable discrete group and $N$ a normal subgroup. If $\mathcal{X}\subseteq\mathrm{Sub}(\Gamma/N)$ is a closed and conjugation-invariant set of subgroups of $\Gamma/N$, then
$J_{\Gamma,q_{N}^{*}(\mathcal{X})}=Q_{N}^{-1}(J_{\Gamma/N,\mathcal{X}}).$
Moreover, if $q^{*}_{N}(\mathcal{X})\in\mathcal{D}_{\Gamma}$, then $\mathcal{X}\in\mathcal{D}_{\Gamma/N}$. Consequently, if $\Gamma$ satisfies the Density Property, then so does $\Gamma/N$.
Proof.
For $X\in\mathcal{X}$, it is easy to see that $\lambda_{\Gamma/q_{N}^{-1}(X)}=\lambda_{(\Gamma/N)/\mathcal{X}}\circ Q_{N}$ and therefore $\ker(\lambda_{\Gamma/q_{N}^{-1}(X)})=Q_{N}^{-1}(\ker(\lambda_{(\Gamma/N)/\mathcal{X}}))$. Hence,
$J_{\Gamma,q^{*}_{N}(\mathcal{X})}=\bigcap_{X\in\mathcal{X}}\ker(\lambda_{\Gamma/q_{N}^{-1}(X)})=Q^{-1}_{N}\big(\bigcap_{X\in\mathcal{X}}\ker(\lambda_{(\Gamma/N)/X})\big)=Q^{-1}_{N}(J_{\Gamma/N,\mathcal{X}}).$
Now, suppose $q_{N}^{*}(\mathcal{X})\in\mathcal{D}_{\Gamma}$. By Lemma 3.3, we may choose a net of positive contractions $(u_{\lambda})\subseteq J_{\Gamma,q^{*}_{N}(\mathcal{X})}\cap\mathbb{C}[\Gamma]$ converging strongly to the identity in $J_{\Gamma,q^{*}_{N}(\mathcal{X})}$. Since $J_{\Gamma,q_{N}^{*}(\mathcal{X})}=Q_{N}^{-1}(J_{\Gamma/N,\mathcal{X}})$, $Q_{N}$ surjects $J_{\Gamma,q_{N}^{*}(\mathcal{X})}$ onto $J_{\Gamma/N,\mathcal{X}}$ and therefore $(Q_{N}(u_{\lambda}))\subseteq J_{\Gamma/N,\mathcal{X}}\cap\mathbb{C}[\Gamma/N]$ is a net of positive contractions converging strongly to the identity in $J_{\Gamma/N,\mathcal{X}}$. By Lemma 3.3, it follows that $\mathcal{X}\in\mathcal{D}_{\Gamma/N}$. ∎
We will now begin to prove the following result.
Theorem 3.6.
The following lemma is probably well-known to experts in group theory - it allows us to approximate an element $a\in C^{*}_{r}(\Gamma)$ by its restriction to finitely many cosets of a normal subgroup $N$. If $\Gamma$ is an amenable discrete group and $N$ a normal subgroup, for $gN\in\Gamma/N$, and $a\in C^{*}_{r}(\Gamma)$, define $a_{gN}\coloneq\delta_{g}\cdot E_{N}(\delta_{g^{-1}}\cdot a)\in C^{*}_{r}(\Gamma)$. Whenever $a_{gN}=0$ for all but finitely many cosets $gN$, the (finite) sum $\sum_{gN}a_{gN}=a$.
Lemma 3.7.
Let $\Gamma$ be an amenable discrete group and $N$ a normal subgroup. Choose a Følner net $(F_{i})$ in $\Gamma/N$ and define $\varphi_{i}:\Gamma/N\to\mathbb{C}$ via $gN\mapsto\frac{|gF_{i}\cap F_{i}|}{|F_{i}|}$. For any $a\in C^{*}_{r}(\Gamma)$ define
$a_{i}\coloneq\sum_{gN\in\Gamma/N}\varphi_{i}(gN)a_{gN}.$
Then, $\|a_{i}\|\leq\|a\|$ for all $i$ and the net $(a_{i})$ converges to $a$.
Proof.
[…]
The fact that $(F_{i})$ is a Følner net implies $(\tilde{\varphi}_{i}(g))$ converges to $1$ for all $g\in\Gamma$. Therefore, $(m_{i}(a))$ converges to $a$ whenever $a\in\mathbb{C}[\Gamma]$, and hence for all $a\in C^{*}_{r}(\Gamma)$ (by boundedness of $m_{i})$. ∎
Proof of Theorem 3.6.
Let $N$ be the (normal) subgroup generated by all $X\in\mathcal{X}$. It suffices to prove the theorem in the case $\Lambda=N$, since $N$ is independent of the ambient group.
Conversely, assume that $J_{N,\mathcal{X}}\cap\mathbb{C}[N]$ is dense in $J_{N,\mathcal{X}}$. By Lemma 3.3 there exists a net of contractions $(u_{beta})\subseteq J_{N,\mathcal{X}}\cap\mathbb{C}[N]$ converging strongly to the identity in $J_{N,\mathcal{X}}$. Since $J_{\Gamma,\mathcal{X}}\cap I_{N}\big(C^{*}_{r}(N)\big)=I_{N}\big(J_{N,\mathcal{X}}\big)$ by Lemma 3.1 (i), we can regard $(u_{\beta})$ as a net in $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$.
Let $a\in J_{\Gamma,\mathcal{X}}$ and fix a Følner net $(F_{i})$ for $\Gamma/N$, with $(a_{i})$ as in Lemma 3.7. By Lemma 3.1, $a_{gN}\in J_{\Gamma,\mathcal{X}}$ for all $gN\in\Gamma/N$. Since $a_{i}$ is a finite linear combination of the $a_{gN}$, it follows that $a_{i}\in J_{\Gamma,\mathcal{X}}$ for all $i$. Moreover, since $\delta_{g^{-1}}\cdot a_{gN}\in J_{N,\mathcal{X}}$, we have $\lim_{\beta}\|a_{gN}\cdot u_{\beta}-a_{gN}\|=0$ for all $gN\in\Gamma/N$ and therefore $\lim_{\beta}\|a_{i}\cdot u_{\beta}-a_{i}\|=0$ for all $i$. Since $(a_{i})$ converges to $a$, $\|a_{i}\|\leq\|a\|$ and $\|u_{\beta}\|\leq 1$ for all $i,\beta$, it follows from the above that $\lim_{\beta}\|a\cdot u_{\beta}-a\|=0$. Lemma 3.3 then implies $J_{\Gamma,\mathcal{X}}\cap\mathbb{C}[\Gamma]$ is dense in $J_{\Gamma,\mathcal{X}}$, completing the proof. ∎
Corollary 3.8.
If $\Gamma$ is an amenable discrete group and $\mathcal{X}\subseteq\mathrm{Sub}(\Gamma)$ is a finite collection of normal subgroups, then $\mathcal{X}\in\mathcal{D}_{\Gamma}$.
Proof.
By Proposition 3.4, we only need to establish the case where $\mathcal{X}=\{N\}$. By Theorem 3.6, it then suffices to establish the case where $\Gamma=N$. In this case, $J_{\Gamma,\{\Gamma\}}$ is the kernel of the trivial representation $\epsilon:C^{*}_{r}(\Gamma)\to\mathbb{C}$. Let $(F_{i})$ be a Følner net for $\Gamma$ and set $v_{i}=\frac{1}{|F_{i}|}\sum_{g\in F_{i}}\delta_{g}$. For any $g\in\Gamma$, we have $\lim_{i}\|\delta_{g}\cdot v_{i}-v_{i}\|=0$, hence (by linearity and density) $\lim_{i}\|a\cdot v_{i}-\epsilon(a)v_{i}\|=0$ for all $a\in C^{*}_{r}(\Gamma)$. By Kesten’s criteria ([5, Theorem 2.6.8(8)]) we have $\|v_{i}\|=1$ for all $i$, so the above limit implies $\lim_{i}\|a\cdot v_{i}\|=|\epsilon(a)|$ for all $a\in C^{*}_{r}(\Gamma)$. Set $u_{i}=\frac{1}{2}(\delta_{e}-v_{i})$ for all $i$. Then $(u_{i})$ is a net of contractions in $J_{\Gamma,\{\Gamma\}}\cap\mathbb{C}[\Gamma]$, and $\lim_{i}\|a\cdot u_{i}-a\|=\lim_{i}\frac{1}{2}\|a\cdot v_{i}\|=\frac{1}{2}|\epsilon(a)|=0$ for any $a\in J_{\Gamma,\{\Gamma\}}$. It follows from Lemma 3.3 that $J_{\Gamma,\{\Gamma\}}\cap\mathbb{C}[\Gamma]$ is dense in $J_{\Gamma,\{\Gamma\}}$, as desired. ∎
Theorem 3.6 shows that in order to determine whether a closed and conjugation-invariant set $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ belongs to $\mathcal{D}_{\Gamma}$, it suffices to assume that $\Gamma$ is generated by the subgroups $X\in\mathcal{X}$. The next result shows that we can also assume $\mathcal{X}$ is minimal in a certain sense.
For any discrete group $\Gamma$ and $\mathcal{X}\subseteq\text{Sub}(\Gamma)$ closed conjugation-invariant, let $\mathcal{X}_{\text{min}}$ be the collection of $X\in\mathcal{X}$ such that $X^{\prime}\subseteq X$ and $X^{\prime}\in\mathcal{X}$ implies $X=X^{\prime}$. Then, $\mathcal{X}_{\text{min}}$ is conjugation-invariant but not necessarily closed.
Theorem 3.9.
Let $\Gamma$ be a discrete group and $\mathcal{X}\subseteq\mathrm{Sub}(\Gamma)$ a closed and conjugation-invariant set of subgroups. Then, for every $X\in\mathcal{X}$, there is $Y\in\mathcal{X}_{\mathrm{min}}$ such that $Y\subseteq X$. Consequently, for $\Gamma$ amenable, we have $J_{\Gamma,\mathcal{X}}=J_{\Gamma,\overline{\mathcal{X}_{\mathrm{min}}}}$ and $\mathcal{X}\in\mathcal{D}_{\Gamma}$ if and only if $\overline{\mathcal{X}_{\mathrm{min}}}\in\mathcal{D}_{\Gamma}$.
Proof.
Fix $X\in\mathcal{X}$. Let $\mathcal{Z}_{X}=\{Y\in\mathcal{X}:Y\subseteq X\}$ and (partial) order $\mathcal{Z}_{X}$ by inclusion. Let $\mathcal{C}\subseteq\mathcal{Z}_{X}$ be a chain in $\mathcal{Z}_{X}$. Then, $\mathcal{C}$ is directed, so we can view it as a net, ordered by inclusion. By compactness of $\mathcal{X}$ the net $\mathcal{C}$ has a convergent subnet $(Y_{\beta})$, and let $Y_{*}$ be its limit point. As $(Y_{\beta})$ is a subnet, for every $Y^{\prime}\in\mathcal{C}$, there is $\beta_{0}$ such that $Y_{\beta_{0}}\subseteq Y^{\prime}$. Therefore, for all $\beta\geq\beta_{0}$, we have $Y_{\beta}\subseteq Y^{\prime}$ and hence $Y_{*}\subseteq Y^{\prime}$. Hence, $Y_{*}\subseteq Y^{\prime}\subseteq X$ for all $Y^{\prime}\in\mathcal{C}$. Therefore, every chain in $\mathcal{Z}_{X}$ has a minimal element. Zorn’s lemma implies $\mathcal{Z}_{X}$ contains a minimal element $Y\in\mathcal{\mathcal{X}}$.
If $X_{1}\subseteq X_{2}$ are subgroups of $\Gamma$, then $\ker(\lambda_{\Gamma/X_{1}})\subseteq\ker(\lambda_{\Gamma/X_{2}})$ (see [2, Appendix E and F]). Therefore, $J_{\Gamma,\mathcal{X}}=\bigcap_{X\in\mathcal{X}}\ker(\lambda_{\Gamma/X})=\bigcap_{X\in\mathcal{X}_{\text{min}}}\ker(\lambda_{\Gamma/X})$. By [3, Proposition 3.3], we have
$\bigcap_{X\in\mathcal{X}_{\text{min}}}\ker(\lambda_{\Gamma/X})=\bigcap_{X\in\overline{\mathcal{X}_{\text{min}}}}\ker(\lambda_{\Gamma/X})$, which completes the proof. ∎
We will now show that abelian groups satisfy the Density Property, and that collections of finite subgroups always belong to $\mathcal{D}_{\Gamma}$. First, we establish a key lemma that unifies the two cases.
Lemma 3.10.
Let $\Gamma$ be an amenable discrete group and $\mathcal{X}\subseteq\mathrm{Sub}(\Gamma)$ a closed and conjugation-invariant set of subgroups. If every $X\in\mathcal{X}_{\mathrm{min}}$ is finitely generated, then $\mathcal{X}_{\mathrm{min}}$ is finite.
Proof.
For each $X\in\mathcal{X}_{\text{min}}$, the set $\mathcal{Z}(X)=\{X^{\prime}\in\mathcal{X}:X\subseteq X^{\prime}\}=\bigcap^{n}_{i=1}\{X^{\prime}\in\mathcal{X}:x_{i}\in X^{\prime}\}$, where $\{x_{1},...,x_{n}\}$ is a finite generating set for $X$. Therefore, $\mathcal{Z}(X)$ is open in $\mathcal{X}$ with respect to the Chabauty topology. Since $\mathcal{Z}(X)\cap\mathcal{X}_{\text{min}}=\{X\}$ for all $X\in\mathcal{X}_{\text{min}}$, the set $\mathcal{X}_{\text{min}}$ is discrete in the compact space $\mathcal{X}$, and is therefore finite. ∎
Now, we see our result for abelian groups and collections of finite subgroups as a simple consequence of Theorems 3.6, Theorem 3.9, Corollary 3.8 and Lemma 3.10.
@@
    $\Gamma$ is countable and abelian, or
@@
    every $X\in\mathcal{X}$ is finite,
@@
Let $\Gamma$ be a countable abelian group. By Proposition 3.2, it suffices to assume that $\Gamma$ is finitely generated. Then, every subgroup of $\Gamma$ is also finitely generated, so Lemma 3.10 implies that $\mathcal{X}_{\text{min}}$ is finite. Since $\Gamma$ is abelian, each $X\in\mathcal{X}_{\text{min}}$ is normal, and hence Corollary 3.8 implies $\mathcal{X}_{\text{min}}\in\mathcal{D}_{\Gamma}$. Theorem 3.9 then implies $\mathcal{X}\in\mathcal{D}_{\Gamma}$.
In the case where every $X\in\mathcal{X}$ is finite, we have that $\mathcal{X}_{\text{min}}$ is finite (by Lemma 3.10) and each $X\in\mathcal{X}_{\text{min}}$ is finite. It follows from [16, Lemma 7.9] (see also the proof of [4, Theorem 4.7]) that the normal subgroup $N$ generated by all $X\in\mathcal{X}_{\text{min}}$ is finite. Since $C^{*}_{r}(N)=\mathbb{C}[N]$, it holds trivially that $\mathcal{X}_{\text{min}}\in\mathcal{D}_{N}$, so Theorems 3.6 and 3.9 imply $\mathcal{X}\in\mathcal{D}_{\Gamma}.$
The following Corollary is immediate by Theorems 2.9 and 3.11.
@@
Let $G$ be an amenable and second-countable étale groupoid, and let $\tilde{G}$ denote the Hausdorff cover groupoid. Assume for every $x\in G^{(0)}$ one of the following holds.
@@
    The isotropy group $G_{x}^{x}$ is abelian.
@@
    The subgroups $X\in\mathcal{X}(x)$ are finite.
@@
Note that condition (II) holds whenever all elements of $\tilde{G}^{(0)}_{\rm{ess}}$ are finite. The class of groupoids described in Corollary 3.12 (II) contains all amenable and second-countable étale groupoids with finite isotropy groups, and all groupoids of contracting self-similar groups (see [4, Corollary 7.13]). More generally, condition (II) is satisfied whenever the unit space of $G$ has finite source and range fibres. Therefore, within the class of amenable and second-countable étale groupoids, the class described in Corollary 3.12 (II) is larger than the class described in [4, Theorem 4.7].
4. Ample Groupoids
The main result of this section is that the algebraic singular ideal $J_{\mathbb{C}}$ is dense in $J\cap\mathscr{C}_{c}(G)$ for any ample groupoid. This reduces the question of whether $J_{\mathbb{C}}$ is dense in $J$ to the corresponding question for $J\cap\mathscr{C}_{c}(G)$. In order to prove this result we will need two preliminary lemmas.

### v2

@@ 4.3. Applications
Let $\Gamma$ be a discrete group and $\mathcal{X}\subseteq\mathrm{Sub}(\Gamma)$ a closed and conjugation-invariant set of subgroups. If every $X\in\mathcal{X}_{\mathrm{min}}$ is finitely generated, then $\mathcal{X}_{\mathrm{min}}$ is finite.
@@
For each $X\in\mathcal{X}_{\text{min}}$, note that $\mathcal{Z}(X)\coloneq\{X^{\prime}\in\mathcal{X}\colon X\subseteq X^{\prime}\}=\bigcap^{n}_{i=1}\{X^{\prime}\in\mathcal{X}\colon x_{i}\in X^{\prime}\}$, where $\{x_{1},...,x_{n}\}$ is a finite generating set for $X$. Therefore, $\mathcal{Z}(X)$ is open in $\mathcal{X}$ with respect to the Chabauty topology, and hence $\{\mathcal{Z}(X)\colon X\in\mathcal{X}_{\text{min}}\}$ forms an open cover of $\mathcal{X}$ by Theorem 4.7. Since $\mathcal{Z}(X)\cap\mathcal{X}_{\text{min}}=\{X\}$ for each $X\in\mathcal{X}_{\text{min}}$, the cover $\{\mathcal{Z}(X)\colon X\in\mathcal{X}_{\text{min}}\}$ admits no proper subcover, so compactness of $\mathcal{X}$ ensures that $\mathcal{X}_{\text{min}}$ is finite. ∎
Recall that a group $\Gamma$ is nilpotent if there is a finite sequence $\{e\}=\Lambda_{0}\trianglelefteq\Lambda_{1}\trianglelefteq\cdots\trianglelefteq\Lambda_{n}=\Gamma$ of subgroups that are normal in $\Gamma$ such that $\Lambda_{i+1}/\Lambda_{i}$ is contained in the centre $Z(\Gamma/\Lambda_{i})$ of $\Gamma/\Lambda_{i}$ for all $i=0,\dots,n-1$. We say $\Gamma$ is finite-by-nilpotent if there is a finite normal subgroup $N\trianglelefteq\Gamma$ such that $\Gamma/N$ is nilpotent.
A subgroup $X\subseteq\Gamma$ is called subnormal if there is a sequence $X=\Lambda_{0}\trianglelefteq\Lambda_{1}\trianglelefteq\cdots\trianglelefteq\Lambda_{k}=\Gamma$ such that $\Lambda_{i}$ is normal in $\Lambda_{i+1}$ for all $i<k$. Such a sequence is called a subnormal series. In a nilpotent group, every subgroup is subnormal. Therefore, in a finite-by-nilpotent group, every subgroup $X$ is contained as a finite index subgroup inside some subnormal subgroup $\Lambda$ ([24, Section 6.3]).
We record a known lemma about finite-by-nilpotent groups that immediately justifies the consideration of Lemma 4.14.
Lemma 4.15.
Let $\Gamma$ be a finitely generated finite-by-nilpotent group. Then every subgroup of $\Gamma$ is finitely generated.
@@
Every finitely generated finite-by-nilpotent group is finite-by-polycyclic, and the property in the lemma holds for such groups ([17]). ∎
We now see our result for finite-by-nilpotent groups and collections of finite (or co-finite) subgroups as a consequence of all our reduction and permanence properties (Theorem 4.5, Theorem 4.7, Theorem 4.9, Corollary 4.6, Proposition 4.8) together with the two previous lemmas.
@@
    every $X\in\mathcal{X}$ is finite,
@@
    every $X\in\mathcal{X}$ is co-finite, or
  3. (III)
    $\Gamma$ is countable and finite-by-nilpotent.
@@
(I) Assume that every $X\in\mathcal{X}$ is finite. By Lemma 4.14 the set $\mathcal{X}_{\text{min}}$ is finite. Since each subgroup $X\in\mathcal{X}_{\text{min}}$ is finite, it follows from [19, Lemma 7.9] (see also the proof of [5, Theorem 4.7]) that the normal subgroup $N$ generated by all $X\in\mathcal{X}_{\text{min}}$ is finite. Then $C^{*}_{r}(N)=\mathbb{C}[N]$, so it holds trivially that $\mathcal{X}_{\text{min}}\in\mathcal{D}_{N}$, and Theorems 4.5 and 4.7 imply $\mathcal{X}\in\mathcal{D}_{\Gamma}$.
(II) Assume every $X\in\mathcal{X}$ is co-finite. Note that if $\Lambda\subseteq\Gamma$ is a subgroup, then $\Lambda\cap X$ is co-finite in $\Lambda$, for every $\Lambda\cap X\in\Lambda\cap\mathcal{X}$. Therefore, Proposition 4.10 allows us to reduce to the special case where $\Gamma$ is finitely generated. Since every co-finite subgroup of a finitely generated group is finitely generated, Lemma 4.14 implies $\mathcal{X}_{\text{min}}$ is finite. Hence, the normal subgroup $N\coloneq\bigcap_{X\in\mathcal{X}_{\text{min}}}X$ is co-finite. Write $\mathcal{Y}=\{X/N\colon X\in\mathcal{X}_{\text{min}}\}$. Then, $\mathcal{Y}$ is a conjugation-invariant set of subgroups in the finite group $\Gamma/N$ with $q^{*}_{N}(\mathcal{Y})=\mathcal{X}_{\text{min}}$. Since $\mathcal{Y}\in\mathcal{D}_{\Gamma/N}$ (automatically due to finiteness of $\Gamma/N$), we have $\mathcal{X}_{\text{min}}=q^{*}_{N}(\mathcal{Y})\in\mathcal{D}_{\Gamma}$ by Theorem 4.9, and hence $\mathcal{X}\in\mathcal{D}_{\Gamma}$ by Theorem 4.7.
(III) Let $\Gamma$ be a countable and finite-by-nilpotent group. Our goal is to show that $\mathcal{X}\in\mathcal{D}_{\Gamma}$. Every subgroup of $\Gamma$ is finite-by-nilpotent so, by Proposition 4.10, we can assume that $\Gamma$ is finitely generated. Then, by Lemma 4.15, every subgroup of $\Gamma$ is finitely generated, so Lemma 4.14 implies that $\mathcal{X}_{\text{min}}$ is finite. By Theorem 4.7 it therefore suffices to assume that $\mathcal{X}$ is finite.
We show that $\mathcal{X}\in\mathcal{D}_{\Gamma}$ by induction on $|\mathcal{X}|$. The base case $|\mathcal{X}|=1$ is covered by Corollary 4.6. Suppose that $|\mathcal{X}|\geq 2$, and assume that whenever a conjugation-invariant collection of subgroups in a finite-by-nilpotent group $\Lambda$ has cardinality less than $|\mathcal{X}|$, then it belongs to $\mathcal{D}_{\Lambda}$. Let $N\subseteq\Gamma$ be the subgroup generated by all elements of $\mathcal{X}$. If $N$ contains every element of $\mathcal{X}$ as a finite index subgroup, then $\mathcal{X}\in\mathcal{D}_{N}$ by (II) above, and hence $\mathcal{X}\in\mathcal{D}_{\Gamma}$ by Theorem 4.5. Suppose otherwise, and choose $X_{0}\in\mathcal{X}$ with infinite index in $N$. Since $\Gamma$ is finite-by-nilpotent, there exists a subnormal subgroup $\Lambda\subseteq\Gamma$ containing $X_{0}$ as a finite-index subgroup. Note that $N\nsubseteq\Lambda$, by assumption. Let $\Gamma=\Lambda_{0}\trianglerighteq\Lambda_{1}\trianglerighteq\dots\trianglerighteq\Lambda_{n}=\Lambda$ be a subnormal series, and let $k\in\{1,\dots,n\}$ be minimal such that $N\nsubseteq\Lambda_{k}$. Define non-empty subsets $\mathcal{X}_{1}\coloneq\{X\in\mathcal{X}\colon X\subseteq\Lambda_{k}\}$ and $\mathcal{X}_{2}=\mathcal{X}\setminus\mathcal{X}_{1}$. Since $\Lambda_{k}$ is normal in $\Lambda_{k-1}$, both $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ are conjugation-invariant collections of subgroups in the finite-by-nilpotent group $\Lambda_{k-1}$ with cardinality strictly less than $|\mathcal{X}|$. By the inductive hypothesis, both $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ belong to $\mathcal{D}_{\Lambda_{k-1}}$. By Proposition 4.8 we have $\mathcal{X}=\mathcal{X}_{1}\cup\mathcal{X}_{2}\in\mathcal{D}_{\Lambda_{k-1}}$, and hence $\mathcal{X}\in\mathcal{D}_{\Gamma}$ by Theorem 4.5. This completes the proof of the theorem. ∎
The following corollary is immediate by Theorems 3.4, Theorem 4.16 and Corollary 4.13, where $\mathcal{X}(x)$ is as defined in Definition 1.3.
@@
Let $G$ be an amenable and second-countable étale groupoid. For each $x\in G^{(0)}$ assume that one of the following holds.
@@
    The isotropy group $G_{x}^{x}$ is finite-by-nilpotent.
@@
    The subgroups $X\in\mathcal{X}(x)$ are all finite.
  3. (III)
    The subgroups $X\in\mathcal{X}(x)$ are all co-finite in $G^{x}_{x}$.
  4. (IV)
    The subgroups $X\in\mathcal{X}(x)$ are all normal in $G^{x}_{x}$.
@@
Note that condition (II) holds whenever all elements of $\tilde{G}^{(0)}_{\rm{ess}}$ are finite. Therefore, within the class of amenable and second-countable étale groupoids, the class described in Corollary 4.17 (II) is larger than the class described in [5, Theorem 4.7].
The final result of this section considers étale group bundles – that is, étale groupoids whose range and source maps are equal.
Corollary 4.18.
Let $G$ be an amenable and second-countable étale group bundle. Then $J\cap\mathscr{C}_{c}(G)$ is dense in $J$.
Proof.
We show that for every $x\in G^{(0)}$, the subgroups $X\in\mathcal{X}(x)$ are all normal in $G^{x}_{x}$. The result then follows by Corollary 4.17 (IV).
Let $x\in G^{(0)}$, $X\in\mathcal{X}(x)$ and choose a net $(x_{\alpha})\subset G^{(0)}$ whose set of limit points is equal to $X$. Take $g\in G^{x}_{x}$ and let $U\subseteq G$ be an open bisection containing $g$. The $x_{\alpha}$ eventually belong to $s(U)$, so there exist $g_{\alpha}\in U$ with $s(g_{\alpha})=x_{\alpha}$. Since $s|_{U}\colon U\to s(U)$ is a homoeomorphism, the net $(g_{\alpha})$ converges to $g$. Then, since $g_{\alpha}=g_{\alpha}x_{\alpha}$ for all $\alpha$, continuity of multiplication on the groupoid ensures that $(g_{\alpha})$ has set of limit points equal to $gX$. However, $r(g_{\alpha})=s(g_{\alpha})$, so we have $g_{\alpha}=x_{\alpha}g_{\alpha}$ for all $\alpha$, and hence continuity of multiplication also ensures that $(g_{\alpha})$ has set of limit points equal to $Xg$. Therefore $gX=Xg$. The element $g\in G^{x}_{x}$ was arbitrary, so $X$ is a normal subgroup of $G^{x}_{x}$, as required. ∎
5. Structure of $J\cap\mathscr{C}_{c}(G)$
In this section we provide explicit instructions for building elements of $J\cap\mathscr{C}_{c}(G)$, and show that functions constructed in this way have dense linear span in $J\cap\mathscr{C}_{c}(G)$. Therefore, when $J\cap\mathscr{C}_{c}(G)$ is dense in the singular ideal $J$, we are able to describe an explicit family of functions in $\mathscr{C}_{c}(G)$ whose linear span is dense in $J$.
Let us describe the approach. One starts with a unit $x\in G^{(0)}$ and a finitely supported function $b\colon G^{x}_{x}\to\mathbb{C}$ that satisfies the linear equations (0.1). Given $h\in G_{x}$, translate the support of $b$ by left-multiplication, and choose open bisections $V_{1},\dots,V_{n}$ containing these translates. Lastly, choose an open neighbourhood $W\subseteq\bigcap_{i=1}^{n}s(V_{i})$ of $x$ such that the open bisections $U_{i}\coloneq V_{i}\cap s^{-1}(W)$ satisfy the conditions of Lemma 5.2. Then, for any $\psi\in C_{c}(W)$, the function $f^{\psi}$ (defined in (5.2)) belongs to $J\cap\mathscr{C}_{c}(G)$ (see Proposition 5.3). Overall, the function $f^{\psi}$ is determined by data in the tuple $(x,h,b,U_{1},\dots,U_{n},\psi)$. By ranging over all valid tuples, one obtains a family of functions whose linear span is dense in $J\cap\mathscr{C}_{c}(G)$ (see Theorem 5.5).
We remark that since $\mathcal{X}(x)$ is invariant under conjugation, an element $b\in\mathbb{C}[G^{x}_{x}]$ satisfies (0.1) if and only if, for every subgroup $X\in\mathcal{X}(x)$, $b$ belongs to the kernel of the associated quasi-regular representation $\lambda_{G^{x}_{x}/X}$.
5.1. Preliminary lemmas
We will need three preliminary results.